OpenStudy (anonymous):

The first equation of a linear system is 8x + 13y = 166. Write a second equation to form a linear system with exactly one solution .

OpenStudy (anonymous):

For the form of Ax +By = C, the slope will be -(B/A). The same equation with a negative inverse slope will be perpendicular and therefore have only one intersection (solution). Just switch the sign of the coefficient of x OR of y and you will get an answer.

OpenStudy (anonymous):

Actually the slope is -(A/B) not -(B/A) but the solution is still the same.

OpenStudy (anonymous):

would -40 + 13y = -830 be the answer? or 8x + 13y = -830

OpenStudy (anonymous):

Actually any slope that ISN'T -(8/13) will work (parallel lines don't touch and have the same slope). Only the coefficients matter for slope, so -40x +13y = -803 would give a single solution.